Consider an insulating sphere with radius a = 9 cm. A charge of -13.3 μC is uniformly distributed throughout this sphere. It is surrounded by a conducting shell. The charge on the inner surface of the shell is q2 and the charge on the outer surface of the shell is q3. The total charge q on the shell is 66.3 μC. Find the charges q2 and q3.
Answer #1

Here,   q2 = 13.3 + q

            q3 =   - 13.3 + q

=>    2q = 66.3

=>   q = 33.15 μC

=> q2 = 13.3 + 33.15 = 46.45 μC

=> q3 =   - 13.3 + 33.15   = 19.85 μC